Given the cost function is

TC = 6L + 3K

Find out the optimal quantities of the two factor using Lagrangian method, if it is given that output is equal to 13.46 = L3/4 . K1/4.



1
Expert's answer
2021-05-04T07:19:06-0400

Given,

"TC=6L+3K"

"13.46=L^{\frac{3}{4}}K^{\frac{1}{4}}"

Using Lagrangian method

L*= minimise cost subject to output constraint

"L^*=6L+3K-\lambda[L^{\frac{3}{4}}K^{\frac{1}{4}}-13.46]"


"\frac{\delta L^*}{\delta L}=6-\lambda[\frac{3}{4}L^{\frac{-1}{4}}K^{\frac{1}{4}}]=0"


"\frac{6(4)}{3L^{\frac{-1}{4}}K^{\frac{1}{4}}}=\lambda"


"8(\frac{L}{K})^{\frac{1}{4}}=\lambda............................................................(1)"


"\frac{\delta L}{\delta K}=3-\lambda[\frac{1}{4}L^{\frac{3}{4}}K^{\frac{1}{4}}]=0"


"\frac{3(4)}{L^{\frac{3}{4}}K^{\frac{-3}{4}}}=\lambda"


"12(\frac{K}{L})^{\frac{3}{4}}=\lambda"


equating

"\lambda'^s, (1) \& (2)"


"8(\frac{L}{K})^{\frac{1}{4}}=12(\frac{K}{L})^{\frac{3}{4}}"


"8L=12K"


"L=\frac{6}{4}K=\frac{3}{2}K"


"L=1.5K"